# GracklePaper (title pending)

### v1

@@ 4.2 Accuracy Test
The One Zone test shows excellent correspondence between Grackle and the fiducial model over the entire time domain for a single region. However, it does not test the model for a large range of initial densities and temperatures. For this, we use a different test. This test is a direct comparison of 106 samples, evolved using both Grackle and the fiducial model. An ideal ML model would have perfect one-to-one agreement between its predictions and Grackle’s calculations. A direct comparison is given in Figure 4, where the calculated values by Grackle are shown on the x-axis, while the predicted values by the fiducial model are shown on the y-axis. In the ideal scenario, all points would lie on a diagonal line with slope one in Figure 4. We can see that for the vast majority of cells, the ML prediction matches the calculations by Grackle, however there is still some scatter present.
@@ 4.2.1 Relative Error
where $x_{G}$ is the value calculated by Grackle, and $x_{i}$ the value predicted by ML. Both values are in logarithmic mass density units, as shown in Table 2. This error can be calculated for each individual species, as well as for the temperature. Since this measure can be quantified quickly for each model, we calculate the relative error for both the fiducial model and the other models considered (i.e. for the Deep, Wide and WideDeep models). This allows for a direct comparison of the performance of each of these models. The distribution of the relative errors for neutral Hydrogen is shown in Figure 5. The distributions of the other species and the temperature can be found in Appendix A. From these figures, we find that the Wide model has the lowest relative error, outperforming the fiducial model. The Deep and WideDeep models both perform worse than the fiducial model, having on average higher relative errors. One possible reason for this is the identical training for the models. Models with more parameters take more training to converge, and the model performance lines up with the number of parameters in each model. Each model received an equal amount of training, for which the fiducial model has converged. It is possible that the Deep and WideDeep models may perform better after a larger amount of training. We will investigate this in a future study.
@@ 4.3 Time Comparison
To compare the scaling of the two methods, several sets of new cells are generated, with different sizes ranging from $10^{3}$ cells to $10^{6}$ cells. We measure the time to completion for each of these sets for Grackle and ML, and repeat this process 20 times to find an average speed, each time with newly initialized cells. Computation is done on a single CPU core for consistency. This disadvantages the ML method, which is optimized for GPU computations. However, most hydrodynamical codes run exclusively on CPUs, so using this hardware gives a more practical comparison between the methods. Future optimisations may include porting chemistry solver calculations to the accelerators on chips.
The resulting computational time and speedups compared to Grackle as a function of the number of cells are shown in Figure 6. The Wide model shows the best speedup, which is expected since it features the fewest amount of layers. The Deep and WideDeep models show worse speedups, with WideDeep being the slowest model. This shows that the possible advantage of the BranchSplit output strategy over the Independent strategy appears smaller than the negative effect of the use of a wider network on computational time. Notable is that that none of the models outperform Grackle for $10^{3}$ cells. This implies that while the ML models scale better to higher cell numbers, Grackle is more lightweight and is a better fit for smaller cell counts.
@@
Therefore, we run the same test, but with different cell generation. Non-equilibrium situations play a more important role in high density regions (Richings, Schaye & Oppenheimer, 2014; Prole et al., 2024). Therefore, a different set of initial ranges is used, with focus on high densities (i.e. star forming regions). The exact values used are shown in Table 3. The results from this test are shown in Figure 7. Comparing these results to those shown in Figure 6, we see a speed-up of the Wide model of a factor of up to 8 compared to Grackle. A potential ideal application of this phenomenon would use a bespoke algorithm to detect non-equilibrium cells from their initial densities and temperatures. This algorithm could potentially also be used to create a hybrid method between Grackle and ML, where Grackle handles the equilibrium cells, where it clearly outperforms ML, while ML tackles the non-equilibrium cells that Grackle spends a significant time on.
@@ 4.4 Iteration
To achieve this we select a cell with a small amount of variance, and use a constant time step of $1$ yr, up until a final time of $10^{3}$ yr, making for $10^{3}$ iterations. The resulting evolution is shown in Figure 8. The top panel shows the evolution over the iterations, while the bottom panel shows the residual error between the ML prediction and Grackle. It’s clear from this plot that the ML method does not align with Grackle for a repeated number of iterations. The discrepancy between the models worsens with the number of iterations, as is shown by the residuals pictured on the bottom plot of Figure 8. This implies that the ML model is not stable over repeated iterations. One cause for this mismatch is that the output space of Grackle is greater than the input space for ML. Grackle, after evolving a cell, may calculate that one or several species densities have values outside of the ranges shown in Table 2. The ML model learns this behaviour appropriately. The next iteration uses the outputs of the previous iteration as input, and since these values lie outside of the ranges shown in Table 2, the ML model is forced to extrapolate. This significantly decreases the quality of the predictions. However, the initial cell used in Figure 8 was chosen such that the values predicted by Grackle would fall within the ranges of the training data shown in Table 2.
@@
The DeepONet architecture only trains on a single timestep, and the minor errors between Grackle and ML predictions compound with repeated application. This is a known issue for several neural network architectures (Stiasny, Misyris & Chatzivasileiadis, 2021), and methods to alleviate this issue for other high-level architectures already exist (e.g Geneva & Zabaras, 2020; List et al., 2024) but these have not been applied to DeepONet in particular. Future work in integrating ML methods into hydrodynamical simulations will have to take this effect into account. Emulating a system of rate equations to work iteratively is however, outside the scope of this work.
@@ 5 Discussion and Conclusion
We demonstrate that our ML models are able to match Grackle to an excellent degree of accuracy (see Figure 5), and provide a computational speedup of up to a factor of 8 for certain density/energy ranges. Our relative errors between the ML result and the results from Grackle average in the range of $10^{-3}$, while Branca & Pallottini (2024) find relative errors around $10^{-2}$. However, they note an approximate speedup of 128, much greater than our results. Their comparison is however against Krome (Grassi et al., 2014), with radiation effects taken into account. This increase in model complexity makes a direct comparison of the speedup challenging, and may also drive the difference in relative error. Nonetheless, our main results agree with those of Branca & Pallottini (2024) in that we both find excellent agreement between the ML approach and the direct chemistry solver for the one zone test and we both clearly show significant computational gains.
@@ 6 Acknowledgements
Appendix A Relative Error Plots
Figure 9: Probability distributions of relative errors between the four ML models and Grackle for the species not shown in Figure 5, as well as the overall density and temperature. These values are calculated using equation 4, based on Grackle and ML predictions of $10^{6}$ samples.

### v2

@@ 4.2 Accuracy Test
The One Zone test shows excellent correspondence between Grackle and the fiducial model over the entire time domain for a single region. However, it does not test the model for a large range of initial densities and temperatures. For this, we use a different test. This test is a direct comparison of 106 samples, evolved using both Grackle and the fiducial model. An ideal ML model would have perfect one-to-one agreement between its predictions and Grackle’s calculations. A direct comparison is given in Figure 4, where the calculated values by Grackle are shown on the x-axis, while the predicted values by the fiducial model for neutral Hydrogen are shown on the y-axis. In the ideal scenario, all points would lie on a diagonal line with slope one in Figure 4. We can see that for the vast majority of cells, the ML prediction matches the calculations by Grackle, however there is still some scatter present. A direct comparison for the other species and the temperature is shown in Appendix A. For these species, Grackle and the fiducial model similarly agree, with some species experiencing more scatter than others. The total density is matched the greatest, showing that the lack of total density evolution by Grackle is matched by the fiducial model. The strongest scatter is present in the temperature, but this is still relatively minor.
@@ 4.2.1 Relative Error
where $x_{G}$ is the value calculated by Grackle, and $x_{i}$ the value predicted by ML. Both values are in logarithmic mass density units, as shown in Table 2. This error can be calculated for each individual species, as well as for the temperature. Since this measure can be quantified quickly for each model, we calculate the relative error for both the fiducial model and the other models considered (i.e. for the Deep, Wide and WideDeep models). This allows for a direct comparison of the performance of each of these models. The distribution of the relative errors for neutral Hydrogen is shown in Figure 5. The distributions of the other species and the temperature can be found in Appendix B. From these figures, we find that the Wide model has the lowest relative error over most of the species, in general matching or improving on the fiducial model, only performing worse to a significant degree on the internal energy. The Deep model matches the performance of the fiducial models for the internal energy and several of the species, however for species such as HI and HeI, the Deep model the performance is worse than either the fiducial or Wide model. The WideDeep model consistently performs worse than the fiducial model, as well as the Wide and Deep models. One possible reason for this is the identical training for the models. Models with more parameters take more training to converge, and the model performance lines up with the number of parameters in each model. Each model received an equal amount of training, for which every model has converged. This was chosen due to computational resources, and to provide a fair comparison between models. While all models have converged, it is possible that an increase in training may lead to further performance gains. We will investigate this in a future study.
@@ 4.3 Time Comparison
To compare the scaling of the two methods, several sets of new cells are generated, with different sizes ranging from $10^{3}$ cells to $10^{6}$ cells. We measure the time to completion for each of these sets for Grackle and ML, and repeat this process 20 times to find an average speed, each time with newly initialized cells. We only measure the time to completion of the call functions of Grackle and the ML models, minimizing potential overheads. Computation is done on a single CPU core for consistency. This disadvantages the ML method, which is optimized for GPU computations. However, most hydrodynamical codes run exclusively on CPUs, so using this hardware gives a more practical comparison between the methods. Future optimisations may include porting chemistry solver calculations to the accelerators on chips. ML calculations are relatively easy to port over to GPUs. This can further increase speedup through parallelization. A traditional solver has to be redesigned for this hardware from scratch, with developments already underway (e.g. Balos et al., 2024).
The resulting computational time and speedups compared to Grackle as a function of the number of cells are shown in Figure 6. The Wide model shows the best speedup, which is expected since it features the fewest amount of layers. The Deep and WideDeep models show worse speedups, with WideDeep being the slowest model. This shows that the possible advantage of the BranchSplit output strategy over the Independent strategy appears smaller than the negative effect of the use of a wider network on computational time. Notable is that none of the models outperform Grackle for $10^{3}$ cells. This implies that while the ML models scale better to higher cell numbers, Grackle is more lightweight and is a better fit for smaller cell counts.
@@
Therefore, we run the same test, but with different cell generation. Non-equilibrium situations play a more important role in high density regions (Richings, Schaye & Oppenheimer, 2014b; Prole et al., 2024). Therefore, a different set of initial ranges is used, with focus on high densities (i.e. star forming regions). The exact values used are shown in Table 3. The results from this test are shown in Figure 7. Comparing these results to those shown in Figure 6, we see a speed-up of the Wide model of a factor of up to 8 compared to Grackle. We note that these non-equilibrium cells, as shown in §4.1 do generally produce less accurate results. This tradeoff will have to be taken into account. A potential application of this phenomenon would use a bespoke algorithm to detect non-equilibrium cells from their initial densities and temperatures. This algorithm could potentially also be used to create a hybrid method between Grackle and ML, where Grackle handles the equilibrium cells, where it clearly outperforms ML, while ML tackles the non-equilibrium cells that Grackle spends a significant time on.
@@ 4.4 Iteration
To achieve this we select a cell with a small amount of variance, and use a constant time step of $1$ yr, up until a final time of $10^{3}$ yr, making for $10^{3}$ iterations. The resulting evolution is shown in Figure 8. The top panel shows the evolution over the iterations, while the bottom panel shows the residual error between the ML prediction and Grackle. It’s clear from this plot that the ML method does not align with Grackle for a repeated number of iterations. The discrepancy between the models worsens with the number of iterations, as is shown by the residuals pictured on the bottom plot of Figure 8. The total density is not constant, failing a first test on convergence. This implies that the ML model is not stable over repeated iterations. One cause for this mismatch is that the output space of Grackle is greater than the input space for ML. Grackle, after evolving a cell, may calculate that one or several species densities have values outside of the ranges shown in Table 2. The ML model learns this behaviour appropriately. The next iteration uses the outputs of the previous iteration as input, and since these values lie outside of the ranges shown in Table 2, the ML model is forced to extrapolate. This significantly decreases the quality of the predictions. However, the initial cell used in Figure 8 was chosen such that the values predicted by Grackle would fall within the ranges of the training data shown in Table 2.
@@
One potential way to constrain the erroneous behaviour of the ML results may be to introduce charge and baryon conservation laws into the initial conditions and after prediction. These may constrain the resulting values, decreasing the rate of divergence. Future work will take this effect into account. However, there is also a more fundamental issue at play in this case. The DeepONet architecture only trains on a single time step, and the minor errors between Grackle and ML predictions compound with repeated application. This is a known issue for several neural network architectures (Stiasny, Misyris & Chatzivasileiadis, 2021), and methods to alleviate this issue for other high-level architectures already exist (e.g Geneva & Zabaras, 2020; List et al., 2024) but these have not been applied to DeepONet in particular. Future work in integrating ML methods into hydrodynamical simulations will have to take this effect into account. Emulating a system of rate equations to work iteratively is however, outside the scope of this work.
@@ 5 Discussion and Conclusion
We demonstrate that our ML models are able to match Grackle to an excellent degree of accuracy (see Figure 5), and provide a computational speedup of up to a factor of 8 for certain density/energy ranges. Our relative errors between the ML result and the results from Grackle average in the range of $10^{-3}$, while Branca & Pallottini (2024) find relative errors around $10^{-2}$. However, they note an approximate speedup of 128, much greater than our results. Their comparison is however against Krome (Grassi et al., 2014), with radiation effects taken into account. This increases the number of chemical reactions, increasing the complexity of the chemical model. Our simpler chemical network takes fewer computations until completion, making Grackle in this configuration faster than the network used by Branca & Pallottini (2024). This increase in model complexity makes a direct comparison of the speedup challenging, and may also drive the difference in relative error. Nonetheless, our main results agree with those of Branca & Pallottini (2024) in that we both find excellent agreement between the ML approach and the direct chemistry solver for the one zone test and we both clearly show significant computational gains.
@@ 6 Acknowledgements
Appendix A Accuracy Plots
Figure 9: Probability distributions of values calculated with Grackle compared to predictions by the fiducial model for the species not shown in Figure 4, as well as the total density and temperature. A diagonal line represents an ideal match between Grackle and the fiducial model.
Appendix B Relative Error Plots
Figure 10: Probability distributions of relative errors between the four ML models and Grackle for the species not shown in Figure 5, as well as the total density and internal energy. These values are calculated using equation 6, based on Grackle and ML predictions of $10^{6}$ samples.